OpenStudy (anonymous):

solve for the variable -19+m=27

OpenStudy (anonymous):

m=46

OpenStudy (anonymous):

Add 19 to both sides.

OpenStudy (anonymous):

since we want to isolate m (get m all by its self) we need to add 19 to both sides and we will get m=46

OpenStudy (anonymous):

-19 + m = 27 +19 +19 m = 27+19 27+19 = 46 so m = 46

OpenStudy (anonymous):

so.... it became... m=46

OpenStudy (anonymous):

okay thanks